simplify:14/-2
5(-4)
7-(-3)
-7(4)
x-5=-3
4x=-8
x/2=-8
-2x=6
x/-5=4
: simplify:14/-2
5(-4)
7-(-3)
-7(4)
x-5=-3
4x=-8
x/2=-8
-2x=6
x/-5=4

Answer by checkley77(3394)

14/-2=-7
5(-4)=-20
7-(-3)=21
-7(4)=-28
x-5=-3 or x=5-3 or x=2
4x=-8 or x=-8/4 or x=-2
x/2=-8 or x=2*-8 or x=-16
-2x=6 or x=6/-2 or x=-3
x/-5=4 or x=-5*4 or x=-20